Operations Report – April 14th

Crew 192 Operations Report 12Apr2018


Name of person filing report: Victoria Varone

Non-nominal systems: Internet. WiFi was down for most of the day, every member of the crew was getting a “WiFi Connected, but No Internet” message. It’s holding now after being reset but we will keep monitoring it.

Generator (hours run): 12 hours

Solar ‒ SOC 84%

Diesel – 75% fuel

Propane – 66% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – Shannon told us we were not to worry about this gas. We have no knowledge of the gas or that 5 gallon container.

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 850 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133477.4

Toilet tank emptied: No. Last emptied April 11, 2018.

ATV’s Used: Red ATV, Blue ATVs #1, #2, and #2

Oil Added? No but checked and OK.

ATV Fuel Used: 20%

# Hours the ATVs were Used today: 6

Notes on ATVs: All working well and ready for use.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 31
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: Yes
Hours: 25.9
Beginning charge: 100%
Ending charge: 91%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Crew is doing well and having a great time going on EVAs and cooking together.

Summary of internet: 697MB at 2040.

Summary of suits and radios: Suits all charging, 3 one-piece suits and 2 two-piece suits used. Radios working fine.

Summary of Hab operations: Two successful EVAs, inspected and tested Musk Observatory (first day of acceptable weather), lots of cooking.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: None

Summary of RAM operations: The trailer water tank remains in the RAM for additional cleaning.

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: Spacesuit 9’s ventilation fan is not working. It worked during our first EVA but is not charging.

Copyright © The Mars Society. All rights reserved. | Main Site